1.The region R is the region bounded by the functions y=x-3 and x=1+y^2. find the volume of the solid obtained by rotating the region R about the y axis. Please include a graph.

2.Find the volume of the solid obtained by rotating the region bounded by the graphs of y=x and y=sqrt(x) about the line x=2. Please include a graph
Answer #1

ax ts = X-3 and 9=x-3 x-3 (5,2 +3 2 2 -0 2 - +りCy-2) 2 sing washe method R 3 R,= 1y Radus oOtes Cuole Radies inner Ccte disc8 mlsinttin o2 2 Volome V +8 2 2 6 2 pa1)-IGh-) TT (3) (9)3(3) S(3) l1 Cobic Unds 55 Solid V= Volumeand Vaboot (Using Shel Method: Rachos of thelR 2 kight o he 8helf -axis k 3- ghell Senfors aSrea R-2 272-)(-) 3/2 S-2TRH= to

Please rate if helpful. Comment if you have any doubt.